# Modal Verification Patterns for Systems

### v2

@@
Abstract
Although they differ in the functionality they offer, low-level systems exhibit certain patterns of design and utilization of computing resources.
In this paper we examine how modalities have emerged as a common structure in formal verification of low-level software, and explain how many recent examples naturally share common structure in the relationship between the modalities and software features they are used to reason about. We explain how the concept of a resource context (a class of system resources to reason about) naturally corresponds to families of modal operators indexed by system data, and how this naturally leads to using modal assertions to describe resource elements (data in the relevant context). We also describe extensions of this idea to ongoing work in formal verification of filesystems code.
1 Introduction
Although they differ in the functionality they offer, low-level systems exhibit certain patterns of design and utilization of computing resources. Systems software, in general, interfaces with an underlying computing substrate such that any software system at any higher level in the software stack can (at least indirectly) utilize the resources of the machine. The last layer of software before the hardware is naturally critical to the correctness of an overall system, as essentially all software built on top of it assumes its correctness. And because hardware is complex and highly diverse, the implementation of those lowest layers of the software stack is typically intricate and naturally error-prone, despite how critical its correctness is. Typically systems software has, as a primary focus, the task of abstracting from hardware details to simplify the construction of higher layers of the stack.
Low-level systems software exhibits certain patterns in their designs, especially when interacting with computing resources. Exploiting certain patterns while designing software has been an important field of study. This survey examines these common patterns in systems software, and their relationships to modalities. We believe that certain properties of modalities enable us to understand and address the verification challenges of these systems by tailoring specification and proof to the same patterns used to design the systems.
Contributions
We argue that modal abstractions can be used to identify and abstract system verification challenges. We justify our perspective by discussing prior systems that have successfully used modalities for system verification, arguing that they fit into the verification design pattern we articulate, and explaining how this approach might apply to other systems’ verification challenges.
@@ 1 Introduction
  • Introducing Modal Resource Contexts: Then we discuss the concept of resource which has already been an essential concept in the design of systems. Inspired by the concept of resource in the systems, we define what a resource and its context are in the modal abstractions.
  • Introducing Nominals for System Resources: Existing logics for systems software must frequently identify individual resources within a given context — nominalization (in the sense of hybrid logic) enables identifying a resource in its context. For example, a filesystem transaction is a context of resources for in-memory updated disk blocks which have not yet been saved to disk. The transaction identifier is used to associate a transaction with a disk-block to be persisted so that, in case of a crash while persisting updated disk-blocks, the filesystem can rollback the already persisted disk-blocks of the transaction, and reach to the previous consistent disk state. We discuss other examples to contextualize the existing systems surveyed.
  • Taxonomy of Current Modal Approaches in System Verification: Based on these concepts defined, we summarize contemporary verification efforts using modal abstractions. We choose them from different domains, for example, reasoning about weak memory models and storage persistence  [5, 6, 7, 47, 13, 14, 9, 33, 48] because we would like to justify that our definitions are not domain dependent.
2 Background on Modal Logics for Programs, and Low-Level Systems Software
This section briefly recalls related concepts from modal logic of particular relevance to our taxonomy, which should be familiar but benefit from fixing notation (Section 2.1), and enough general background on the systems software concepts organized by our taxonomy for the connections to modalities to be clear (Section 2.2).
2.1 Hybrid Logic, Dynamic Logic, and Nominals in Program Logics
@@
We believe that for the kinds of informal reasoning and the sorts of data structures discussed in the previous section, ideas from modal logic are a promising approach to formalization and verification. Broadly speaking modal logics incorporate modal operators, which take as arguments a proposition expected to be true in another time [39], place [19, 18, 1, 17, 35], circumstance or point-of-view [25, 23], and result in a proposition true in the current time, place, or circumstance at which the truth of the use of the modal operator is being evaluated. Classic examples include modal necessity $\square P$ describing that $P$ is necessarily true, $\mathsf{G}~{}P$ meaning $P$ is true globally (i.e., forever from this time onwards in temporal logics), or $K_{i}(P)$ describing that a particular participant $i$ knows that $P$ is true. The latter is an example of multimodal logic, where there is an family of modalities (modal operators) parameterized by some dimension of interest (there, participants).
Dynamic logic [40], a (multi)modal logic of weakest preconditions [12] which works with modalities of the form $[p](P)$, which states that in the current program state, if program $p$ is run then afterwards $P$ will hold (modulo non-termination). This same idea is used to encode Hoare triples in the Iris program logic [28], using the same encoding as in Pratt’s original presentation, where a Hoare triple $\{P\}C\{Q\}$ is encoded as $P\rightarrow[C](Q)$. Iris is now one of the most successful foundational program logics, and its rich support for extending the base logic with additional modalities [31] means that modal logics, and in particular dynamic logic, now play an outsized role in program verification — many challenging proofs of program correctness use a substructural dynamic logic containing many additional modalities which interact with the weakest precondition modality.
@@
Hybrid logics [2, 18, 17, 1, 4] are a class of modal logics that add two primary new concepts to a logic. Nominals ł$\l\in\textsf{Loc}$ uniquely identify points in a model (e.g., a particular state), so there is exactly one point in the model where a nominal Loc is true. Satisfaction operators are modal operators indexed by nominals, which enable claims about the truth of another proposition at some arbitrary point in the model identified by a nominal. Traditionally @$_{l}\varphi$ asserts that $\varphi$ is true in the state identified by $l$. Conceptually, given our context, we might like $l$ to indicate truth in a particular data structure $l$. However, most prior work on hybrid logics works in classical logics rather than a substructural setting like separation logic. Newer work surveyed here adapts this idea to deal with different structures specified in separation logic.
@@
An example use of state naming explicitly in the assertion appears in program logics with usage protocols (e.g. state transition systems resembling typestate [46, 16], which ensure the operations applied to some data occur in a specified order, and that invariants for various conceptual states of the data — such as a file being open or closed — are respected). Protocol assertions are annotated with the name of the last (abstract) state at which the protocol is ensured. A classic application of this idea to systems is Halpern et al.’s work adapting modal logics of knowledge to deal with distributed systems [24, 22, 21]. In most of that line of work, $\mathcal{K}_{a}(P)$ indicates that the node $a$ in the system knows or possesses knowledge of $P$ (for example, a Raft node may “know” a lower bound on the commit index). Alternatively, a modality @$_{i}(P)$ may represent that $P$ is true of/at the specific node $i$ [19] (e.g., that node $i$ has stored a certain piece of data to reliable storage). These permit capturing specific concepts relevant to the correctness (and reasoning about correctness) of a certain class of systems, involving facts about specific named entities in the system.
In each of these cases, the fact that these facts are described using modalities with the core modal property $P\Rightarrow Q$, then $M(P)\Rightarrow M(Q)$ means, for example, that if a process $p$ knows that the commit index is greater than 5 (e.g., $\mathcal{K}_{p}(\mathsf{commitIndex}>5)$) no extra work is required to conclude that the node knows it is greater than 3 (i.e., that $\mathcal{K}_{p}(\mathsf{commitIndex}>3)$), because this follows from standard properties of modal operators as described above. In contrast, if verification instead used a custom assertion $\mathsf{minCommitIndex}(5)$ to represent the former knowledge, one would need to separately provide custom reasoning to conclude $\mathsf{minCommitIndex}(3)$. As our survey shows, the value of the modal view (which grants both intuitive specifications and natural reasoning principles) is becoming apparent in systems software verification.
2.2 Virtualization in Systems Software
@@
Operating System (OS) kernels virtualize memory locations and quantity (via virtual memory and paging [11]). Distributed language runtimes may virtualize addresses, even when processes may migrate across machines [27]. Filesystems virtualize locations on disk [43, 26, 44, 3]. Programs built on top of the corresponding systems software layer work logically at the level of these virtualized resources, and it makes sense to specify the systems software directly in terms of those abstractions.
@@
Accessing virtualized resources via translation is a common way to virtualize notions of location (e.g., virtual memory addresses, inodes or object IDs instead of disk addresses [3, 26, 44, 43]). B-trees, page tables, and related structures both behave like maps, when the corresponding physical resources exist as such just in a different location. Control over the lookup process (e.g, handling the case of a missing translation entry) allows for additional flexibility, such as filling holes in sparse files, or demand paging (both from disk or lazily populating anonymous initially-zeroed mappings). Although the realization of these maps may differ from system to system based on the context (and sometimes hardware details), they are semantically — logically — partial maps, worth treating as such in verification.
2.2.1 Virtualization of Memory Locations
A typical general purpose computer virtualizes memory resources in RAM. A program asking for a memory unit from an operating system kernel is served with a memory address that is virtualized such that it may be mapped to a physical RAM location or not. As our informal definition suggests, the common technique for virtualization of memory locations in OS kernels relies translating (mapping) virtual addresses into physical ones. The conceptual address-translation map is implemented with processor’s page-table trees as shown in Figure 1. In this Figure, we see a typical 4 levels of page tables, a virtual address on the top indexing into the different levels of the page-table tree (the larger rectangles along the bottom are blocks of memory which serve as nodes of the tree along a particular lookup path). Fabrication of virtual addresses out of limited physical addresses is provided with entries in the page-table tree where each points to a physical address aligned to 4KB (4096 byte) boundaries. An address translation requires traversal of a series of tables starting from level 4, and it is conventionally referred as a page-table walk. The traversal ends in the level 1 page table with the final lookup to the actual page of physical memory holding the requested data, and the low-order 12 bits being used to index into this page.
Figure 1: x86-64 page table lookups.
@@ 2.2.1 Virtualization of Memory Locations
2.2.2 Virtualized Resources in Filesystems
@@
Another important computer resource computer resource that is virtualized for access is disk blocks. Filesystems are software components that abstract disk access on behalf of normal programs Similar to how VMMs using page-table trees to fabricate more memory addresses than what actually exists, filesystems uses indexing maps for fabricating more virtual disk-page addresses which are conventionally called disk-pages to be distinguished from physical disk-block addresses. When a regular program accesses a file, it requests a specific logical offset from the start of that file’s contents. As with virtual memory, the filesystem may map adjacent logical blocks of a single file to physically disparate parts of the disk (or even to different disks), though this translation mapping is implemented entirely in software (as opposed to the hardware cooperation that exists for memory virtualization).
@@
Filesystems not only handle the address-translation from a disk-page to a disk-block but also need to handle different modes of computing where to find disk resources: a disk block may live on a disk in stable storage, or the most up-to-date version of a block may live only in memory (for the filesystem to write to disk at a later time, batched for efficiency). To handle the consistency of the disk resources in the existence of different modes of execution, filesystems employ different policies when indexing the disk resource. Which data-structure is going to be used to implement address-translation from disk-pages to disk-blocks? What is going to be the update-policy to the indexing map implementing address-translation which defines the consistency-policy for the disk resources? Different filesystem implementations give different answers to these questions.
@@
While exceptions exist [3, 43, 44, 26], most filesystems will continue to store a given block of a given file in the same physical disk block for as long as the file exists. These filesystems with in-place updates need to implement some form of journalling [20, 8] to ensure that a crash partway through an in-place update can be rolled back using an undo journal that records the original contents before in-place modification (so original contents could be restored after a crash) or completed using a write-ahead journal that first writes new data in unused space before also updating in place (so the journal entries could be re-applied after a crash). Because disk blocks associated with a file could now be in many states (on-disk, journaled but not updated in place, or both) the specification of these systems is complex. And because some states are only possible after a system failure, each line of filesystem code has two conceptual post-conditions: one for continued execution, and one for the stable state established in the event of a hardware failure immediately after that operation.
2.2.3 Reference-Counting Memory Reclamation
In a typical configuration, allocating, deallocating and referencing memory locations for client code leaves the memory layout ideally in a form of DAG (directed acyclic graph) with a unique root, which resembles the first stack pointer referencing the allocated memory. DAG abstraction instead of recursive decent tree allows non-duplication on the reference counts for the identical memory objects.

### v3

@@
Abstract.
Although they differ in the functionality they offer, low-level systems exhibit certain patterns of design and utilization of computing resources. In this paper we examine how modalities have emerged as a common structure in formal verification of low-level software, and explain how many recent examples naturally share common structure in the relationship between the modalities and software features they are used to reason about. We explain how the concept of a resource context (a class of system resources to reason about) naturally corresponds to families of modal operators indexed by system data, and how this naturally leads to using modal assertions to describe resource elements (data in the relevant context).
separation logic, modal logic, verification, operating systems
journalyear: 2025copyright: ccconference: 13th Workshop on Programming Languages and Operating Systems; October 13–16, 2025; Seoul, Republic of Koreabooktitle: 13th Workshop on Programming Languages and Operating Systems (PLOS ’25), October 13–16, 2025, Seoul, Republic of Koreadoi: 10.1145/3764860.3768337isbn: 979-8-4007-2225-7/25/10ccs: Theory of computation Modal and temporal logicsccs: Theory of computation Logic and verificationccs: Software and its engineering Operating systems
1. Introduction
Although they differ in the functionality they offer, low-level systems exhibit certain patterns of design and utilization of computing resources. Systems software, in general, interfaces with an underlying computing substrate such that any software system at any higher level in the software stack can (at least indirectly) utilize the resources of the machine.
The last layer of software before the hardware is naturally critical to the correctness of an overall system, as essentially all software built on top of it assumes its correctness. And because hardware is complex and highly diverse, the implementation of those lowest layers of the software stack is typically intricate and naturally error-prone, despite how critical its correctness is. Typically systems software has, as a primary focus, the task of abstracting from hardware details to simplify the construction of higher layers of the stack.
Interacting with computing resources is the essential point which shapes the design of low-level systems software. Exploiting these design choices has been an important field of study. This survey examines these common patterns in systems software, and their relationships to modalities. We believe that certain properties of modalities enable us to understand and address the verification challenges of these systems by tailoring specification and proof to the same patterns used to design the systems.
Contributions We argue that modal abstractions can be used to identify and abstract system verification challenges. We justify our perspective by discussing prior systems that have successfully used modalities for system verification, arguing that they fit into the verification design pattern we articulate, and explaining how this approach might apply to other systems’ verification challenges.
@@ 1. Introduction
  • Introducing Modal Resource Contexts: Then we discuss the concept of resource which has already been an essential concept in the design of systems. Inspired by the concept of resource in systems, we define what a resource and its context are in the modal abstractions. For example, the transaction abstraction, which behaves like a container for in-memory dirty disk pages, can be considered a resource context for any dirty in-memory disk page resource associated with it.
  • Introducing Nominals for System Resources: Existing logics for systems software must frequently identify individual resources within a given context — nominalization (in the sense of hybrid logic) enables identifying a resource in its context. For example, a filesystem transaction is a context of resources for in-memory updated disk blocks which have not yet been saved to disk. The transaction identifier is used to associate a transaction with a disk block to persist so that, in case of a crash while persisting updated disk-blocks, the filesystem can rollback the already persisted disk-blocks of the transaction, and restore the previous consistent disk state. We discuss other examples to contextualize the existing systems surveyed.
  • Taxonomy of Current Modal Approaches in System Verification: Based on these concepts, we summarize contemporary verification efforts using modal abstractions. We choose them from different domains, for example, reasoning about weak memory models and storage persistence  (Chajed, 2022; Chajed et al., 2023, 2019; Vindum et al., 2025; Doko and Vafeiadis, 2016, 2017; Dang et al., 2019; Kuru and Gordon, 2025; Wagner et al., 2024; Gordon, 2019) and using different programming logics such as Iris  (Jung et al., 2018) and Dafny  (Leino, 2010) because we would like to justify that our definitions are neither domain nor programming logic dependent.
2. Background on Modal Logics for Programs, and Low-Level Systems Software
This section briefly recalls related concepts from modal logic of particular relevance to our taxonomy (Section 2.1), and enough general background on the systems software concepts organized by our taxonomy for the connections to modalities to be clear (Section 2.2).
2.1. Hybrid Logic, Dynamic Logic, and Nominals in Program Logics
@@
We believe that for the kinds of informal reasoning and the sorts of data structures discussed in the previous section, ideas from modal logic are a promising approach to formalization and verification. Broadly speaking modal logics incorporate modal operators, which take as arguments a proposition expected to be true in another time (Pnueli, 1977), place (Gordon, 2019; Goranko, 1996; Areces et al., 2001; Gargov and Goranko, 1993; Murphy et al., 2004), circumstance or point-of-view (Hintikka, 1962; Halpern and Moses, 1985), and result in a proposition true in the current time, place, or circumstance at which the truth of the modal operator is being evaluated. Classic examples include modal necessity $\square P$ describing that $P$ is necessarily true, $\mathsf{G}(P)$ meaning $P$ is true globally (i.e., forever from this time onwards in temporal logics), or $K_{i}(P)$ describing that a particular participant $i$ knows that $P$ is true. The latter is an example of multimodal logic, where there is an family of modalities (modal operators) parameterized by some dimension of interest (there, participants).
Dynamic logic (Pratt, 1976), a (multi)modal logic variant of weakest preconditions (Dijkstra, 1975; hahnle2022dijkstra; ahrendt2025many) which works with modalities of the form $[p](P)$, which states that in the current program state, if program $p$ is run then afterwards $P$ will hold (modulo non-termination).
@@
Hybrid logics (Blackburn and Seligman, 1995; Goranko, 1996; Gargov and Goranko, 1993; Areces et al., 2001; Braüner, 2010) are a class of modal logics that add two primary new concepts to a logic. Nominals $\ell\in\textsf{Loc}$ uniquely identify points in a model (e.g., a particular state in the state space the logic is used to reason about), so there is exactly one point in the model where a nominal Loc (as a formula) is true. Satisfaction operators are modal operators indexed by nominals, which enable claims about the truth of another proposition at the point in the model identified by a chosen nominal. Traditionally $_{\ell}\varphi$ asserts that $\varphi$ is true in the state identified by $\ell$. Conceptually, given our context, we might like $l$ to indicate truth in a particular data structure $\ell$. However, most prior work on hybrid logics works in classical logics rather than a substructural setting like separation logic. Newer work (Chajed, 2022; Chajed et al., 2023, 2019; Vindum et al., 2025; Doko and Vafeiadis, 2016, 2017; Dang et al., 2019; Kuru and Gordon, 2025; Wagner et al., 2024):w surveyed here adapts this idea to deal with different structures specified in separation logic.
@@
An example use of state naming explicitly in the assertion appears in program logics with usage protocols (e.g. state transition systems resembling typestate (Strom and Yemini, 1986; Garcia et al., 2014), which ensure the operations applied to some data occur in a specified order, and that invariants for various conceptual states of the data — such as a file being open or closed — are respected). Protocol assertions are annotated with the name of the last (abstract) state at which the protocol is ensured. A classic application of this idea to systems is Halpern et al.’s work adapting modal logics of knowledge to deal with distributed systems (Halpern and Moses, 1990; Halpern and Fagin, 1989; Halpern, 2017). In most of that line of work, $\mathcal{K}_{a}(P)$ indicates that the node $a$ in the system knows or possesses knowledge of $P$ (for example, a Raft node may “know” a lower bound on the commit index). Alternatively, a modality $_{i}(P)$ may represent that $P$ is true of/at the specific node $i$ (Gordon, 2019) (e.g., that node $i$ has physically stored a certain piece of data to reliable storage). These permit capturing specific concepts relevant to the correctness (and reasoning about correctness) of a certain class of systems, involving facts about specific named entities in the system.
In each of these cases, the fact that these facts are described using a modality $M$ with the core modal property that if $P\Rightarrow Q$, then $M(P)\Rightarrow M(Q)$. This means, for example, that if a process $p$ knows that the commit index is greater than 5 (e.g., $\mathcal{K}_{p}(\mathsf{commitIndex}>5)$) no extra work is required to conclude that the node knows it is greater than 3 (i.e., that $\mathcal{K}_{p}(\mathsf{commitIndex}>3)$), because this follows from standard properties of modal operators as described above. In contrast, if verification instead used a custom assertion
$\mathsf{minCommitIndex}(5)$
to represent the former knowledge, one would need to separately provide custom reasoning to conclude
$\mathsf{minCommitIndex}(3)$
As our survey shows, the value of the modal view (which grants both intuitive specifications and natural reasoning principles) is becoming apparent in systems software verification.
2.2. Virtualization in Systems Software
@@
Operating System (OS) kernels virtualize memory locations and quantity (via virtual memory and paging (Denning, 1970)). Distributed language runtimes may virtualize addresses, even when processes may migrate across machines (Jul et al., 1988). Filesystems virtualize locations on disk (Rodeh et al., 2013; Hitz et al., 1994; Rosenblum and Ousterhout, 1992; Bonwick et al., 2003). Programs built on top of the corresponding systems software layer work logically at the level of these virtualized resources, and it makes sense to specify the systems software directly in terms of those abstractions.
@@
Accessing virtualized resources via translation is a common way to virtualize notions of location (e.g., virtual memory addresses, inodes or object IDs instead of disk addresses (Bonwick et al., 2003; Hitz et al., 1994; Rosenblum and Ousterhout, 1992; Rodeh et al., 2013)). B-trees, page tables, and related structures both behave like maps, when the corresponding physical resources exist as such just in a different location. Control over the lookup process (e.g, handling the case of a missing translation entry) allows for additional flexibility, such as filling holes in sparse files, or demand paging (both from disk or lazily populating anonymous initially-zeroed mappings). Although the realization of these maps may differ from system to system based on the context (and sometimes hardware details), they are semantically — logically — partial maps, worth treating as such in verification.
2.2.1. Virtualization of Memory Locations
A typical general purpose computer virtualizes memory resources in RAM. A program asking for a memory unit from an operating system kernel is served with a memory address that is virtualized such that it may be mapped to a physical RAM location or not. As our informal definition suggests, the common technique for virtualization of memory locations in OS kernels relies on translating (mapping) virtual addresses into physical ones. The conceptual address translation map is implemented with processor’s page table trees as shown in Figure 1. In this Figure, we see a typical 4 levels of page tables, a virtual address on the top indexing into the different levels of the page table tree (the larger rectangles along the bottom are blocks of memory which serve as nodes of the tree along a particular lookup path). Fabrication of virtual addresses out of limited physical addresses is provided with entries in the page table tree where each points to a physical address aligned to 4KB (4096 byte) boundaries. An address translation requires traversal of a series of tables starting from level 4. The traversal ends in the level 1 page table with the final lookup to the actual page of physical memory holding the requested data, and the low-order 12 bits being used to index into this page.
Figure 1. x86-64 page table lookups.
@@ 2.2.1. Virtualization of Memory Locations
2.2.2. Virtualized Resources in Filesystems
@@
Another important computer resource that is virtualized for access is disk blocks. Filesystems are software components that abstract disk access on behalf of normal programs Similar to how VMMs using page table trees to fabricate more memory addresses than what actually exists, filesystems use indexing maps for fabricating more virtual disk page addresses which are conventionally called disk pages, to be distinguished from physical disk block addresses. When a regular program accesses a file, it requests a specific logical offset from the start of that file’s contents. As with virtual memory, the filesystem may map adjacent logical blocks of a single file to physically disparate parts of the disk (or even to different disks), though this translation mapping is implemented entirely in software (as opposed to the hardware cooperation that exists for memory virtualization).
@@
Filesystems not only handle the address translation from a disk-page to a disk-block but also need to handle different modes of computing where to find disk resources: a disk block may live on a disk in stable storage, or the most up-to-date version of a block may live only in memory (for the filesystem to write to disk at a later time, batched for efficiency). To handle the consistency of the disk resources in different modes of execution, filesystems employ different policies when indexing the disk resource. Which data-structure is going to be used to implement address translation from disk-pages to disk-blocks? What is going to be the update policy to the indexing map implementing address translation, which defines the consistency policy for the disk resources? Different filesystem implementations give different answers to these questions.
@@
While exceptions exist (Bonwick et al., 2003; Rodeh et al., 2013; Rosenblum and Ousterhout, 1992; Hitz et al., 1994), most filesystems will continue to store a given block of a given file in the same physical disk block for as long as the file exists. These filesystems with in-place updates need to implement some form of journalling (Hagmann, 1987; Chutani et al., 1992) to ensure that a crash partway through an in-place update can be rolled back using an undo journal that records the original contents before in-place modification (so original contents could be restored after a crash) or completed using a write-ahead journal that first writes new data in unused space before also updating in place (so the journal entries could be re-applied after a crash). Because disk blocks associated with a file could now be in many states (on-disk, journaled but not updated in place, or both) the specification of these systems is complex. And because some states are only possible after a system failure, each line of filesystem code has two conceptual post-conditions: one referring to the successfully stored disk pages from a disk page container (e.g, a disk tree) , which can be uniquely identified by the last successful commit identifer (e.g., a transaction identifier or snapshot identifier), as the current view of the disk in the specification for continued execution, and one referring to the previous view of the disk in the specification for the stable state established in the event of a hardware failure immediately after that operation.
2.2.3. Reference Counting Memory Reclamation